OpenStudy (anonymous):

subtract and simplify. (6-i)-(7+4i) I got the answer of 37+17i. is that right?

OpenStudy (zzr0ck3r):

distribute the negative sign and what do you have?

OpenStudy (caozeyuan):

no! where does your 37 come form?

OpenStudy (caozeyuan):

open the parenthesis first

OpenStudy (anonymous):

Distribute the negative sign to get 6-i-7-4i. Now simplify

OpenStudy (caozeyuan):

6-i-7-4i=(6-7)-(i+4i)so it's -1-5i

OpenStudy (anonymous):

(6-i)-(7+4i) =6-i-7-4i= -1-5i is the ans

OpenStudy (anonymous):

okay i did the multiplying and then subtracting. but this does make more sense. thank you.

OpenStudy (caozeyuan):

you got to be more careful reading the question

OpenStudy (anonymous):

yes I do. I get in a hurry to get the work done and then miss the key word.